# An Efficient Learning Control Framework With Sim-to-Real for String-Type Artificial Muscle-Driven Robotic Systems

### v2

@@ III-A Robotic Eye
III-B Robotic Wrist
Research into artificial muscle-driven robotic wrist joints holds considerable importance in several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields. We designed a parallel robotic wrist prototype driven by three equally spaced compound TCAs [30] as illustrated in Fig. 3. The base plate remains stationary, while the upper plate, also known as the end effector, moves via three muscle strings. Linkages between the two plates are constructed to support two degree-of-freedom motion of the upper plate [38]. Given this configuration, we developed a simplified dynamics model under the assumption of a fixed rotational center for the upper plate. Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation. This necessitates adjusting the applied actuating voltages across the three muscle strings to drive the upper plate from its initial equilibrium state to predefined target angles.
The reference frame $O_{xyz}$ is fixed on the moving upper plate, with the $z$-axis perpendicular to the upper plate and pointing outward. $\theta,\phi,$ and $\psi$ represent the pitch, roll, and yaw angles, respectively. Similar to the robotic eye, the system state variables are defined as
$\bm{x}=[\theta,\dot{\theta},\phi,\dot{\phi},\Delta T_{1},\Delta T_{2},\Delta T_{3}]$ (9)
where $\dot{\theta}$ and $\dot{\phi}$ represent the corresponding angular velocities, and $\Delta T_{i}$ is the temperature difference of muscle $m_{i}$ with respect to the ambient environment, with $i\in\{1,2,3\}$. The system output vector is represented as $\bm{y}=[\theta,\dot{\theta},\phi,\dot{\phi}]$. Likewise, the muscle’s temperature is not used as an input for the DRL agent, and thus, the state vector is defined as
$s_{t}=[\theta_{t},\dot{\theta}_{t},\phi_{t},\dot{\phi}_{t},\theta^{*},\phi^{*}].$ (10)
Here, $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate. Additionally, the input action vector $a_{t}$ of the system is defined as
$a_{t}=[V_{1},V_{2},V_{3}]$ (11)
where $V_{i}$ represents the actuating voltage applied to muscle $m_{i}$, constrained by $0\leq V_{i}\leq 10$. The output and control weight matrices of the reward funtion are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. When the tracking error is sufficiently small, a bonus reward is granted, i.e.,
IV Learning Control Framework
@@
This section provides a comprehensive overview of the proposed learning control framework, as illustrated in Fig. 4. The soft actor-critic (SAC) serves as the baseline DRL algorithm. To enhance data efficiency, we employ a two-stage training strategy. Initially, a standard PID controller is initialized to generate imperfect trajectories for the DRL agent. Subsequently, the agent interacts directly with the simulation environment to update its policy. All transition data observed from the environment undergoes an state-based augmentation strategy before being stored into the replay buffer. To address the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics at the beginning of each training episode. Eventually, the trained actors (control policies) are able to be directly applied in real-world robotic systems to achieve desired control performance.
IV-A Soft Actor-Critic (SAC)
SAC, a state-of-the-art deep reinforcement learning algorithm, finds extensive application across various tasks [39]. Traditional DRL algorithms primarily maximize cumulative rewards, i.e.,
$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}[\mathcal{R}(s_{t},a_{t})]$ (13)
where $\pi$ represents the policy. However, SAC adopts an entropy-maximizing approach, with its learning objective formulated as
$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}[\mathcal{R}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))].$ (14)
where $\mathcal{H}(\pi(\cdot|s_{t}))$ represents an entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter. This approach enhances the robustness of DRL and accelerates its training process [31].
Figure 5: Schematic diagram of the network architecture. The actor (left) and the critic (right) networks share the same structure but have different parameters. Each is composed of a fully-connected layer with ReLU activation as the input layer, a recurrent GRU [40] unit, and a fully-connected output layer.
In addition, SAC is an off-policy algorithm that leverages the actor-critic architecture. The neural network structures utilized in this paper are illustrated in Fig. 5. To process the sequential information crucial for implementing dynamics randomization and handling partial observability, we employ a Gated Recurrent Unit (GRU) [40]. We select GRU over over commonly-used Long Short-Term Memory (LSTM) due to its efficiency and ease of training [40]. Noise is added to the observation input to the actor. The hyperparameters of the SAC algorithm are detailed in Table I and remain consistent across all ablation studies conducted in this paper.
TABLE I: SAC Hyper-parameters
Parameter Value
Learning rate of actor $1\times 10^{-4}$
Learning rate of critic $1\times 10^{-4}$
Replay buffer size $1\times 10^{5}$
Mini-batch size $20$
Entropy target $-\text{dim}(\mathcal{A})$ (e.g., $-3$ for robotic wrist)
Soft update factor $\tau$ $0.005$
Discount factor $\gamma$ $0.99$
Hidden layer size of GRU $256$
Gradient steps $k$ $40$
Target networks update interval $1$
IV-B Randomization of Muscle Dynamics
The core idea behind dynamics randomization is to introduce randomness into specific dynamic parameters within the simulator during training. This approach allows the DRL agent to learn a policy adaptable to various dynamics conditions, including those encountered in real-world experiments. Widely adopted in diverse robotic applications [41, 42, 43], this method serves as inspiration for our work. Specifically, we randomize several critical parameters within the simulated dynamics environment. Among the robotic systems discussed in this paper, the dynamic parameter set of the muscles, denoted by $\mu$, holds pivotal importance. Therefore, we establish a distribution $\rho_{\mu}$ to encompass the variability of muscle dynamic parameters for dynamics randomization purposes. Additionally, we incorporate Gaussian noise into the observation inputs of the policy to emulate the noise commonly present in real-world sensor data.
Regarding the SCP muscles driving the robotic eye and the TCA muscles driving the parallel robotic wrist, their thermomechanical models are approximately as [29]
$F=k(x-x_{0})+b\dot{x}+c(T(t)-T_{\text{amb}}).$ (15)
Here, $x$ and $x_{0}$ represent the loaded length and unloaded length of the actuator, respectively. $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect of the actuator, respectively. $T(t)$ denotes the current temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment. The thermoelectric model of the actuator is expressed as
$C_{\text{th}}\frac{dT(t)}{dt}=P(t)-\lambda(T(t)-T_{\text{amb}}).$ (16)
Here, $C_{\text{th}}$ represents the thermal mass of the actuator, and $\lambda$ represents the absolute thermal conductivity of the actuator. $P(t)$ stands for the heat power applied to the actuator, calculated as $P(t)=V^{2}/R$, where $V$ is the voltage applied to the actuator and $R$ is the electrical resistance. In summary, the dynamic parameter set for randomization is
$\mu:=\{k,b,c,C_{\text{th}},\lambda,R\}$ (17)
@@ IV-B Randomization of Muscle Dynamics
TABLE II: Dynamics parameters of randomization and their respective range in the robotic eye
@@
$k$ $(N/cm)$ $0.25$ uniform$(0.8,1.2)$
$b$ $(N\cdot s/cm)$ $0.01$ uniform$(0.9,1.1)$
$c$ $(N/^{\circ}C)$ $0.0055$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(Ws/^{\circ}C)$ $0.28$ uniform$(0.8,1.2)$
$\lambda$ $(W/^{\circ}C)$ $0.094$ uniform$(0.85,1.15)$
@@
$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/s)$ normal$(0,0.05)$
Table II and III outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems. $\rho_{\mu}$ is centered on the mean values derived from system identification to align the data distribution in simulation more accurately with that in reality. The range of randomization for each parameter is determined based on the distribution observed across different real muscle strings, as outlined in previous work [29]. During the training process, the physical parameters of the actuators are randomized within the scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the beginning of each training episode and remain constant throughout. Whereas, the observation noise varies at each time step and across different Euler angles. While this paper employs a simplified linear dynamic model for artificial muscles, this method is expected to be readily transferred to more complex muscle dynamics models.
IV-C Bootstrap via PID
On-line DRL agents mainly rely on random exploration [31]. However, due to the slow response of artificial muscles, random exploration policies may struggle to explore the state space comprehensively and rapidly, resulting in low sample efficiency. To address this issue, we draw inspiration from previous work [44, 45] and boost the training process by providing demonstrations into initial replay buffer. While PID controllers demand extensive tuning for optimal control performance, their implementation is relatively straightforward, given that moderate performance suffices. In the case of muscle-driven robotic systems, PID policies typically suffice to approach the target state. Hence, they provide trajectories that explore the defined state space more comprehensively, serving as an advantageous starting point for the DRL learning process. Subsequently, the agent will interact with the environment and continue to refine its policy, thus eventually outperforming PID and achieving desired control performances.
Before the DRL training process begins, we initialize a PID controller that offers a barely acceptable (and typically not excellent) control performance. Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the the rewards. Differing from traditional behavior cloning methods, we don’t utilize a supervised loss during demonstration learning. Consequently, the agent does not strictly follow the PID control policy but leverages the trajectories generated by PID solely to bootstrap the initial training.
IV-D Augmentation of State Vector
Data augmentation refers to a class of techniques aimed at enhancing the size and quality of training datasets, thereby facilitating the development of more robust deep learning models [46]. The fundamental concept of data augmentation involves transforming limited data into diverse, functionally equivalent datasets. Given the labor-intensive and computationally expensive nature of collecting training data in robotic systems, we integrate data augmentation into the agent’s learning process. While prior research [47, 48, 49] has explored data augmentation in reinforcement learning, existing studies have primarily focused on pixel-based RL scenarios, where agents rely on image-based observations. In this paper, we present a straightforward and adaptable approach to state-based data augmentation, better suited and more effective for our targeted robotic systems.
TABLE III: Dynamics parameters of randomization and their respective range in the robotic wrist
@@ IV-D Augmentation of State Vector
$k$ $(N/mm)$ $0.21$ uniform$(0.8,1.2)$
$b$ $(N\cdot s/cm)$ $0.63$ uniform$(0.9,1.1)$
$c$ $(N/^{\circ}C)$ $0.0707$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(Ws/^{\circ}C)$ $3.06$ uniform$(0.8,1.2)$
$\lambda$ $(W/^{\circ}C)$ $0.1189$ uniform$(0.85,1.15)$
@@
$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/s)$ normal$(0,0.05)$
As commonly acknowledged, rewards serve as pivotal guides in the learning process of DRL agents. Therefore, our objective is to generate more data transitions with different rewards. By combining (2) and (3), the reward $r_{t}$ is then calculated as
$r_{t}=\mathcal{R}(\bm{x}_{t},\bm{y}^{*},a_{t}).$ (18)
Here, $\bm{x}_{t}$ and $a_{t}$ are directly related to system dynamics, whereas the target output vector $\bm{y}^{*}$ is manually set and changeable. Consequently, the proposed augmentation of state vector proceeds as follows. At each time step $t$, we fix the motion state vectors $\bm{x}_{t},\bm{x}_{t+1}$ and the action $a_{t}$, and vary the target vector $\bm{y}^{*}$. In this paper, we introduce a random additive term into $\bm{y}^{*}$ as follows
$\bm{y}^{*}_{\text{new}}=\bm{y}^{*}\pm\delta\bm{Z}.$ (19)
$(s_{t},a_{t},r_{t},s_{t+1})\xrightarrow{AUG}(s^{\prime}_{t},a_{t},r^{\prime}_{t},s^{\prime}_{t+1}).$ (20)
Algorithm 1 Augmentation of State-based Data
1:  Generate rollout $\tau=(s_{0},a_{0},r_{0},...,s_{T})$ from environment
2:  for $i=1$ to $n$ do
3:     for each $s_{t},a_{t},r_{t}$ in $\tau$ do
4:        Create a new state vector $s_{t}\rightarrow s^{\prime}_{t}$
5:        Calculate a new reward $r^{\prime}_{t}=\mathcal{R}(s^{\prime}_{t},a_{t})$
6:     end for
7:     Get a new rollout $\tau^{\prime}_{i}=(s^{\prime}_{0},a_{0},r^{\prime}_{0},...,s_{T})$
8:  end for
TABLE IV: Simulation Parameters of Robotic Systems
Description Symbol Value
Robotic Eye Total time duration of one episode $t_{\text{total}}$ $15$s
Action time step $t_{a}$ $0.5$s
Total number of training episodes $N$ $2000$
Random or boost episodes $M$ $250$
Absolute value of $\theta$ bound $\theta_{\text{max}}$ $15^{\circ}$
Absolute value of $\psi$ bound $\psi_{\text{max}}$ $10^{\circ}$
Robotic Wrist Total time duration of one episode $t_{\text{total}}$ $30$s
Action time step $t_{a}$ $0.5$s
Total number of training episodes $N$ $3500$
Random or boost episodes $M$ $500$
Absolute value of $\theta$ bound $\theta_{\text{max}}$ $15^{\circ}$
Absolute value of $\phi$ bound $\phi_{\text{max}}$ $15^{\circ}$
@@
The procedure of using augmentation for a state-based trajectory $\tau$ is provided in Algorithm 1. Each time the agent interacts with the environment and gets a new trajectory, this method creates $n$ different trajectories at a very low computational cost. We set $n=10$ in both robotic eye and wrist systems during training. In this paper, we denote the bootstrap method by the letter ’B’, augmentation method by the letter ’A’, and randomization of muscle dynamics by the letter ’R’. Algorithm 2 provides a detailed overview of the entire training procedure for the SAC algorithm with all three proposed improvements (i.e., SAC-BAR).
V Experimental Results
This section presents the experimental results of the proposed efficient learning control framework within two testing robotic platforms. We first present the training results of the algorithm in simulation environments, validating the effectiveness of the proposed bootstrap and augmentation methods in enhancing training data efficiency. Subsequently, we demonstrate the results of an ablation study conducted on the robotic wrist platform, verifying the effectiveness of dynamics randomization in improving the real-world control performance. Finally, experimental results on robotic eye are provided to demonstrate the generalizability of the proposed control framework across different robotic systems of interest.
(a) Robotic Eye
(b) Robotic Wrist
Figure 6: Reward trajectories with respect to the training episode number in ablation study. Ours refers to the proposed SAC-BAR algorithm, where B, A, and R indicate bootstrap, augmentation, and randomization techniques, respectively. Performance is evaluated in simulation with random dynamics every $10$ training episodes. The shaded area indicates the standard deviation of the mean reward based on running each experiment with three different random seeds.
Figure 7: Experimental trajectories of vanilla SAC algorithm on the robotic wrist under target angles ($\theta^{*}=-5^{\circ},\phi^{*}=7.5^{\circ}$) and ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$), respectively. (a), (c) The system output angles $(\theta,\phi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
Figure 8: Experimental trajectories of our proposed SAC-BAR algorithm on the robotic wrist under target angles ($\theta^{*}=-5^{\circ},\phi^{*}=7.5^{\circ}$) and ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$), respectively. (a), (c) The system output angles $(\theta,\phi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
V-A Training Results
We established simulation environments for two robotic systems to train DRL agents and conduct preliminary tests. Table IV lists the simulation parameters for each robotic system. The overall duration of a single episode $t_{\text{total}}$ and the action time step $t_{a}$ depend on the rise time and response speed constraints inherent to the hardware of the systems. The values of $N$ and $M$ are empirical parameters contingent on the complexity of the systems. For simpler dynamics, smaller values for $N$ and $M$ facilitate the learning process for the DRL agent in acquiring an effective control policy. For algorithms not using the bootstrap method, random exploration is employed in the first $N$ episodes. Target angles are confined within the motion range of the robotic device. With the majority of training time spent on dynamics calculation, the time step for dynamics simulation is set at $0.01$s to balance simulation accuracy and computational efficiency. Training and testing in simulation are conducted on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU. The algorithms are implemented in PyTorch and trained using the Adam optimizer.
In each training episode, a target state is randomly selected from a uniformly distributed range. For instance, in the robotic eye scenario, the target angles adhere to $\theta^{*}\sim\mathcal{U}[-15^{\circ},15^{\circ}]$ and $\psi^{*}\sim\mathcal{U}[-10^{\circ},10^{\circ}]$. The robotic agent endeavors to reach the target from the initial state throughout the episode. For algorithms incorporating dynamics randomization, a new set of muscle dynamics parameters $\mu$ is sampled at the beginning of each episode by drawing values for each parameter from their respective range. Additionally, simulated sensor noise is introduced to the observed state at every action step. In the absence of dynamics randomization, muscle parameters are kept constant at their mean values, and no simulated noise is added during the training process.

### v3

@@ V Learning Control Framework
This section provides a comprehensive overview of the proposed learning control framework, as illustrated in Fig. 5. The soft actor-critic (SAC) algorithm serves as the baseline DRL method. To improve data efficiency, we employ a two-stage training strategy. Initially, a standard PID controller is initialized to generate imperfect trajectories for the DRL agent. Subsequently, the agent interacts directly with the simulation environment to update its policy. All transition data collected from the environment undergoes a state-based augmentation strategy before being stored into the replay buffer. To mitigate the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics at the beginning of each training episode. Eventually, the trained actors (control policies) can be directly implemented in real-world robotic systems to achieve the desired control performance.
V-A Soft Actor-Critic
Soft Actor-Critic (SAC) [29], a state-of-the-art DRL algorithm, is widely adopted as a strong baseline for continuous control tasks in robotics due to its robustness, sample efficiency, and model-free nature. Unlike traditional DRL algorithms, SAC incorporates an entropy-maximizing approach, with its learning objective defined as:
$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}\left[\mathcal{R}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))\right],$ (10)
where $\mathcal{H}(\pi(\cdot|s_{t}))$ represents the entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter. This entropy-regularized objective improves the robustness of DRL and accelerates its training [24]. To process sequential information critical for addressing partial observability issue and implementing dynamics randomization, we integrate a Gated Recurrent Unit (GRU) [41] as the hidden layer in the actor-critic architecture.
V-B Randomization of Muscle Dynamics
@@
TABLE I: Dynamics parameters of randomization and their respective ranges in the robotic eye
@@
[…]
$c$ $(\text{N}/^{\circ}\text{C})$ $0.0055$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $0.28$ uniform$(0.8,1.2)$
$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.094$ uniform$(0.85,1.15)$
@@
$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
The dynamic model of the artificial muscles considered in this paper is based on the literature[26], where the muscle’s elasticity behavior is linearized to establish a dynamical system comprising a mechanical model characterized by stiffness and damping, given by
$F=k(x-x_{0})+b\dot{x}+c(T(t)-T_{\text{amb}}),$ (11)
where $x$ and $x_{0}$ denote the loaded length and unloaded length of the actuator, respectively, while $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect coefficients of the actuator, respectively. $T(t)$ denotes the current actuator temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment. The thermoelectric model of the actuator is expressed as:
$C_{\text{th}}\frac{dT(t)}{dt}=P(t)-\lambda(T(t)-T_{\text{amb}}),$ (12)
where $C_{\text{th}}$ represents the thermal mass of the actuator (measured in $\text{Ws}/^{\circ}\text{C}$), $P(t)$ indicates the heat applied to the actuator, $\lambda$ is the absolute thermal conductivity of the actuator with respect to the ambient environment (in $\text{W}/^{\circ}\text{C}$). The applied power, $P(t)$, is controlled through the voltage applied to the muscle, expressed as $P(t)=V(t)^{2}/R$.
TABLE II: Dynamics parameters of randomization and their respective ranges in the robotic wrist
@@
[…]
$c$ $(\text{N}/^{\circ}\text{C})$ $0.0707$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $3.06$ uniform$(0.8,1.2)$
$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.1189$ uniform$(0.85,1.15)$
@@
$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
The core idea behind dynamics randomization is to introduce randomness into specific dynamic parameters within the simulator during training. This approach allows the DRL agent to learn a policy adaptable to various dynamics conditions, including those encountered in real-world experiments. Specifically, we randomize several critical muscle parameters within the simulated dynamics environment. A distribution $\rho_{\mu}$ of the muscle parameter set $\mu$ is established to encompass the variability of muscle dynamic parameters for dynamics randomization purposes. This approach is expected to be particularly effective in our setup, as artificial muscles are known to exhibit significant variability due to factors such as material degradation, slight manufacturing differences, and environmental temperature changes. According to the muscle models (11) and (12), the dynamic parameter set $\mu$ for randomization is given by
$\mu:=\left\{k,b,c,C_{\text{th}},\lambda,R\right\}.$ (13)
Table I and II outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems. $\rho_{\mu}$ is centered on the mean values derived from system identification to align the data distribution in simulation more accurately with that in reality. To strike the right balance between randomization and fidelity, we carefully tune the range and distribution of randomized parameters based on empirical observations of muscle behavior in real-world scenarios. Specifically, the range of each parameter is determined based on the distribution observed across different real muscle strings, as outlined in [26]. This approach allows us to maintain a high degree of fidelity while simultaneously enabling the controller to generalize effectively to real-world uncertainties. Details on the impact of varying randomization ranges on experimental results are discussed in Section VII-C.
During the training process, the physical parameters of the actuators are randomized within predefined scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the start of each training episode and remain constant throughout. Additionally, Gaussian noise is added to the policy’s observation inputs to simulate the noise typically present in real-world sensor data. This observation noise varies at each time step and across different state variables. Although this study employs a simplified dynamic model for artificial muscles, the proposed method can be extended to more complex muscle dynamics models to further enhance control performance.
V-C Bootstrap via PID
On-line DRL agents primarily rely on random exploration [24]. However, due to the slow response characteristics of artificial muscles, random exploration policies often fail to explore the state space comprehensively and efficiently, leading to low sample efficiency. To address this limitation, we accelerate the early-stage training by incorporating demonstration data into the initial replay buffer. For muscle-driven robotic systems, PID controllers are typically sufficient to approach target states and deliver moderate control performance. As such, they can generate trajectories that explore the defined state space more comprehensively, providing a beneficial starting point for the DRL learning process.
Specifically, we initialize a PID controller that offers a barely acceptable (and typically not excellent) control performance before DRL training begins. Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the rewards. Unlike traditional behavior cloning methods, we do not employ a supervised loss for demonstration learning, allowing the agent to avoid strictly mimicking the PID control policy. Subsequently, the agent continues interacting with the environment to refine its policy, eventually surpassing the PID controller and achieving superior control performance.
V-D Augmentation of State Vector
Data augmentation refers to a class of techniques aimed at increasing the size and improving the quality of training datasets, thereby facilitating the development of more robust deep learning models[42]. The fundamental concept of data augmentation involves transforming limited data into diverse and functionally equivalent datasets. Given the labor-intensive and computationally expensive nature of collecting training data in robotic systems, we integrate data augmentation into the agent’s learning process. This paper proposes a novel and straightforward state-based data augmentation technique designed for our targeted robotic systems.
@@ V-D Augmentation of State Vector
As commonly acknowledged, rewards serve as pivotal guides in the learning process of DRL agents. Therefore, our objective is to generate additional data transitions with different rewards. As described in Section III, the reward $r_{t}$ is calculated as:
$r_{t}=\mathcal{R}(\bm{x}_{t},\bm{y}^{*},a_{t}).$ (14)
Here, $\bm{x}_{t}$ and $a_{t}$ are directly related to system dynamics, whereas the target output vector $\bm{y}^{*}$ is manually set and changeable. Consequently, the proposed augmentation of state vectors proceeds as follows. At each time step $t$, we fix the motion state vectors $\bm{x}_{t},\bm{x}_{t+1}$ and the action $a_{t}$, and vary the target vector $\bm{y}^{*}$. In this paper, we introduce a random additive term into $\bm{y}^{*}$ as follows
$\bm{y}^{*}_{\text{new}}=\bm{y}^{*}\pm\delta\bm{Z}.$ (15)
$(s_{t},a_{t},r_{t},s_{t+1})\xrightarrow{AUG}(s^{\prime}_{t},a_{t},r^{\prime}_{t},s^{\prime}_{t+1}),$ (16)
which creates different transition tuples at a negligible computational cost.
In this paper, we denote the bootstrap method by the letter ’B’, augmentation method by the letter ’A’, and the randomization of muscle dynamics by the letter ’R’. Thus, our proposed SAC algorithm with all three improvements is denoted by SAC-BAR. The pseudo code of our overall SAC-BAR learning control framework is presented in Algorithm 1.
Figure 6: Reward trajectories with respect to the training episode number. (a) Comparison results between our SAC-BAR and SOTA methods including REDQ, DroQ, RLPD, and X-QL. (b) Results of ablation study where ‘B’, ‘A’, and ‘R’ indicate bootstrap, augmentation, and randomization techniques, respectively.
VI Simulation Experiments
In this section, we present the training results from simulation environments, comparing our proposed SAC-BAR against several SOTA reinforcement learning algorithms in terms of sample efficiency on two robotic systems. Additionally, we perform ablation studies to evaluate the effectiveness of the proposed bootstrap and augmentation methods.
VI-A Training Setup
At the beginning of each training episode, a target state is randomly selected from a uniformly distributed range. For instance, in the robotic eye scenario, the target angles adhere to $\theta^{*}\sim\mathcal{U}[-15^{\circ},15^{\circ}]$ and $\psi^{*}\sim\mathcal{U}[-10^{\circ},10^{\circ}]$. The robotic agent strives to reach the target from the initial state throughout the episode. In this paper, simulation environments are all written in Python. Training and testing in simulation are conducted on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU. The algorithms are implemented in PyTorch and trained using the Adam optimizer. The learning rate of DRL algorithms is set to $3\times 10^{-4}$ and the hidden layer size is fixed at $256$. The training hyperparameters $N$ and $M$ are empirically chosen based on the system’s complexity. Specifically, $N=2000$ and $M=250$ are used for the robotic eye, while $N=3500$ and $M=500$ are used for the robotic wrist. When the bootstrap method is not applied, random exploration is employed during the initial $N$ warm-up episodes. These hyperparameters are kept consistent across all comparison and ablation studies.
VI-B Comparison Results Against SOTA Methods
Several SOTA sample-efficient RL algorithms, designed for state-based observations and continuous action spaces, are selected for comparison. REDQ[34] integrates an ensemble of Q-functions with in-target minimization across a random subset of these functions. This allows for training with a high Update-To-Data (UTD) ratio, significantly improving sample efficiency. DroQ[35] equips the Q-functions with dropout connections and layer normalization techniques, maintaining a high UTD ratio. RLPD[36] leverages offline data during online learning by loading demonstrations into the replay buffer and oversampling them during training. X-QL[37] directly models soft optimal values and introduces a novel optimization framework based on the Gumbel distribution.
We trained the proposed SAC-BAR on two robotic systems and compared its performance with these SOTA approaches. Performance was measured by the online per-step average reward achieved by the agents, with each algorithm trained across four independent runs. Fig. 6(a) presents the comparative results between SAC-BAR and the selected SOTA RL methods. In the robotic eye system, SAC-BAR achieves performance comparable to (and slightly better than) DroQ and RLPD, while significantly outperforming REDQ and X-QL in terms of sample efficiency. In the robotic wrist system, SAC-BAR not only significantly outperforms RLPD and DroQ but also demonstrates moderate improvements in convergence speed over REDQ and X-QL. Notably, REDQ’s performance relies on a high UTD ratio and an ensemble of Q-functions, resulting in high computational cost; for instance, REDQ requires 18.2 hours for 3,500 training episodes, whereas SAC-BAR completes the same training in only 4.1 hours under our experimental setup. Moreover, X-QL exhibited significant instability during the early stages of training in both robotic systems. Overall, these results demonstrate that SAC-BAR is better suited than the existing SOTA RL methods for the string-type artificial muscle robots targeted in this study.
VI-C Ablation Study in Simulation
To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, an ablation study was conducted on two robotic systems. Three variants of our proposed framework were tested as follows: (1) w/o A: This variant excludes the state vector augmentation method. (2) w/o B: This variant removes the bootstrap via PID method. (3) w/o AB: This baseline variant excludes both mechanisms. Each algorithm was trained and evaluated in simulation under randomized dynamics across four independent runs.